# Rank jumps and Multisections of elliptic fibrations on K3 surfaces

### v1

@@ Acknowledgements
It is our pleasure to thank Bert van Geemen for reading and commenting on our manuscript, and Wim Nijgh for pointing out the quartic K3 that appears in Example 6.3 and discussing details about it. The second author also would like to thank Brendan Hassett for a discussion on the ramification of trisections on elliptic K3 surfaces and the Institute for Advanced Study for their hospitality and stimulating atmosphere. Part of her work on this material is supported by the NWO XL grant Rational points: new dimensions, the James D. Wolfensohn Fund and the National Science Foundation under Grant No. DMS-1926686. The first author is member of the INDAM-GNSAGA.

### v2

@@ Acknowledgements
It is our pleasure to thank the referee for their helpful comments and for pointing out a mistake in the previous version of the paper, Bert van Geemen for reading and commenting on our manuscript, and Wim Nijgh for pointing out the quartic K3 that appears in Example 6.3 and discussing details about it. The second author also would like to thank Brendan Hassett for a discussion on the ramification of trisections on elliptic K3 surfaces and the Institute for Advanced Study for their hospitality and stimulating atmosphere. Part of her work on this material is supported by the NWO XL grant Rational points: new dimensions, the James D. Wolfensohn Fund and the National Science Foundation under Grant No. DMS-1926686. The first author is member of the INDAM-GNSAGA.